Solve for x: √(x+7)=x-5
Answer by stanbon(75887)

Solve for x: √(x+7)=x-5
-----
Square both sides:
x+7 = x^2-10x+25
---
x^2-11x+18 = 0
x^2-9x-2x+18 = 0
x(x-9)-2(x-9) =0
(x-9)(x-2) = 0
x = 2 or x = 9
-------------------------
Check those "answers":
Only x = 9 is acceptable.
-------------------------
Cheers,
Stan H.